A solution of hydrochloric acid has a concentration of 4.50g L-1 .

 (i) Calculate the hydronium ion concentration in mol L-1 .

(ii) Calculate the pH of the solution.

 


1
Expert's answer
2021-09-30T03:05:13-0400

Solution:

hydrochloric acid = HCl

hydronium ion = H3O+


The molar mass of HCl is 36.458 g/mol

Hence,

(4.50 g HCl L-1) × (1 mol HCl / 36.458 g HCl) = 0.123 mol L-1

C(HCl) = 0.123 mol L-1


HCl(aq) + H2O(l) → H3O+(aq) + Cl-(aq)

HCl is a strong acid.

Hence,

[H3O+] = C(HCl) = 0.123 mol L-1

[H3O+] = 0.123 mol L-1


We can convert between [H3O+] and pH using the following equation:

pH = - log[H3O+]

pH = - log(0.123) = 0.91

pH = 0.91


Answers:

(i): [H3O+] = 0.123 mol L-1

(ii): pH = 0.91